OpenStudy (anonymous):

A line passes through (6, –4) and (–7, –1). Identify the point-slope and slope-intercept form of an equation of the line.

OpenStudy (ghazi):

y-y1= slope (x-x1), slope= -1-(-4)/ -7-6= -3/13 , now using the given points solve for your slope intercept form

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

y=-3/13x+22??

OpenStudy (ghazi):

y+4= -3/13(x-6) this is it

OpenStudy (anonymous):

yes thats the point slope @ghazi but i also need the slope intercept but i dont really know how to do that its wither between y=-3/13x-34/13 or y=-3/13x+22